OpenStudy (anonymous):

What is the equation of the vertical asymptote of the graph of y=log5(x-2)

OpenStudy (lgbasallote):

hint \[\huge \log _a (0) = \infty\] does that help?

OpenStudy (anonymous):

so it is infinity

OpenStudy (lgbasallote):

no...

OpenStudy (lgbasallote):

\[y = \log_a (0) \implies \text{vertical asymptote}\]

OpenStudy (lgbasallote):

let me give a real example..

OpenStudy (lgbasallote):

\[\huge y = \log_2 (3x + 4)\] the verical asymptote of this is when 3x + 4 = 0 so i solve for x 3x + 4 = 0 3x = -4 x = -4/3 does that help?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what is the vertical asymtote for \[\huge y = \log_5 (x-2)\]

OpenStudy (anonymous):

x=2

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

so what is the equation

OpenStudy (lgbasallote):

x = 2

OpenStudy (lgbasallote):

that's the equation

OpenStudy (anonymous):

ok